What is the mean and standard deviation of the sampling distribution of the means. n=4


1
Expert's answer
2022-03-08T13:04:07-0500

"\mu_{\bar X}=\mu."

"\sigma_{bar X}=\frac{\sigma}{\sqrt{4}}=\frac{\sigma}{2}."